OpenStudy (anonymous):

how to solve the proportion 3y-8/12=y/5

OpenStudy (anonymous):

Multiply both sides by 5, then combine the y's and solve.

OpenStudy (haseeb96):

3y-8/12 = y /5 by cross multiplication 5( 3y - 8) = 12y 15y -40 =12y 15y -12y =40 3y=40 y=40/3 y=13.33 is the correct answer

OpenStudy (anonymous):

15 y - 40/12 = y 14 y = 40/12 y = 0.238 Checks 3(13.33) - 0.666 = (13.33)/5 does not check